# Bayesian optimisation of poloidal field coil positions in tokamaks

### v1

@@ Abstract
Abstract. The tokamak is a world-leading concept for producing sustainable energy via magnetically-confined nuclear fusion. Identifying where to position the magnets within a tokamak, specifically the poloidal field (PF) coils, is a design problem which requires balancing a number of competing economic, physical, and engineering objectives and constraints. In this paper, we show that multi-objective Bayesian optimisation (BO), an iterative optimisation technique utilising probabilistic machine learning models, can effectively explore this complex design space and return several optimal PF coil sets. These solutions span the Pareto front, a subset of the objective space that optimally satisfies the specified objective functions. We outline an easy-to-use BO framework and demonstrate that it outperforms alternative optimisation techniques while using significantly fewer computational resources. Our results show that BO is a promising technique for fusion design problems that rely on computationally demanding high-fidelity simulations.
Keywords. Bayesian optimisation $\bullet$ Poloidal field coils $\bullet$ Spherical tokamak $\bullet$ MHD equilibria $\bullet$ FreeGS
UKAEAUnited Kingdom Atomic Energy Authority, Culham Campus, Abingdon, Oxfordshire, OX14 3DB, United Kingdom
UCLCentre for Artificial Intelligence, Department of Computer Science, University College London, London, WC1V 6LJ
1 Introduction
1.1 Motivation and aims
A spherical tokamak is a torus-shaped device with a low aspect ratio that uses strong magnetic fields to confine and control a thermonuclear fusion plasma, with the goal of producing fusion energy (Freidberg, 2008). The Spherical Tokamak for Energy Production (STEP), currently in the design phase and targeting completion in 2040 (Chapman et al., 2024), is one of a few ongoing international fusion research and development projects based on the spherical tokamak concept. To deliver fusion power to the grid on such a short timescale, researchers are increasingly designing next-generation tokamaks in silico with multi-physics simulations, many of which require high-performance computing (HPC) resources.
@@ 1.1 Motivation and aims
Our focus here will be on the design of the poloidal field (PF) coil system, which plays a critical role in controlling the position and shape of the plasma in both the core and divertor regions of a tokamak (Lim et al., 2010). In particular, some coils are crucial for managing the vertical stability of elongated plasmas, such as those in spherical tokamaks, where the higher elongation can lead to larger vertical instability, risking disruption without appropriate control (Anand et al., 2023). By generating poloidal magnetic fields, the PF coil system ensures the plasma remains in equilibrium, balancing the inward-facing magnetic forces produced by the coils against the outward-facing pressure-driven forces generated by the plasma (Wesson and Campbell, 2011, Chp. 3.1). The design of the system—in terms of the coil positions, size, and shape—will have a significant impact on plasma performance and stability and will therefore need to satisfy a number of competing (and often conflicting) constraints. In the plasma, for example, constraints are required to ensure X-points form in specific locations (for stability), strikepoints hit the correct divertor plates (for heat management), and total current density limits on the PF coils are not exceeded. In terms of the tokamak itself, the locations/sizes of the coils will inevitably be constrained by the vacuum vessel, diagnostic systems, and maintenance ports (to name but a few).
In addition to constraints, there will be a number of objectives related to the desired operational plasma conditions that we wish the chosen coil set to minimise or maximise (depending on the objective). This could include minimising the coil size to reduce fabrication, construction, and installation costs or could include minimising current flows to reduce power consumption and structural stresses from forces produced by the coils (Coleman and McIntosh, 2020). Moreover, we may wish to optimise certain properties of the plasma in the divertor chambers in order to minimise heat loads on plasma facing components and improve exhaust performance (Hudoba et al., 2023b). Simultaneously satisfying both the objectives and constraints will require the solution of a complex optimisation problem that needs to be tackled in a systematic, computationally efficient manner.
In this paper, we will perform multi-objective Bayesian optimisation (BO) on an earlier baseline design of the STEP PF coil system (Hudoba et al., 2023a). Our aims are to:
@@
    encourage more widespread adoption of BO for the in silico design of interlinked components on future tokamak devices to save time, minimise financial costs, and improve plasma performance.
@@ 1.2 Related work
Despite these advantages, the adoption and application of BO in fusion engineering and design has, so far, remained relatively limited. Brown et al. (2024b, a) use BO to optimise the current-profiles in STEP. Their goal was to improve six key properties of the safety factor profile. They additionally demonstrate that BO performs better than a genetic algorithm with the same number of black-box function evaluations. Mehta et al. (2024) use BO to find the parameters such as neutral beam injection power, plasma current, and plasma elongation in the DIII-D tokamak that safeguard against disruption during the ramp-down phase. Similarly, Pusztai et al. (2023) use BO to mitigate the impact of disruptions in ITER by exploring how injected deuterium and neon can minimise runaway electron currents, transported heat, and quench time post-disruption. Järvinen et al. (2022) also investigate runaway electron currents using BO as an advanced sampling method to help calibrate uncertainty and minimise the discrepancy between simulations and experimental data. For fusion component design, Humphrey et al. (2023) demonstrates BO to minimise stresses in parametrised divertor monoblock under fusion conditions. The most relevant work to ours is that of Nunn et al. (2023), who use multi-objective BO to optimise TF coil shapes to reduce both financial costs and magnetic ripples (which affect plasma stability and performance). In contrast, our approach deals with more computationally expensive, failure-prone plasma equilibrium simulations without analytic objective/constraint functions, necessitating the use of a classifier alongside the surrogate model (as we will see later on).
@@
The work here is inspired by that of Hudoba et al. (2023a), in which the authors seek to optimise the STEP PF coil system by minimising deviations of key plasma parameters (from a baseline scenario which we adopt) and coil currents, while maximising divertor performance metrics. Using a free-boundary equilibrium solver, thousands of potential PF coil sets are sampled and evaluated (in a Monte Carlo-type approach) before optimal solution sets are identified heuristically. We aim to provide and fully outline an alternative, much more data efficient, framework for carrying out similar multi-objective optimisation that can return a STEP equilibrium similar to the baseline.
@@
There are also a number of areas in fusion design where BO has yet to be applied but could potentially offer significant benefits. For example, parameter scans for optimal magnetic sensor placement, as explored for TCV (Romero and Svensson, 2013) and SPARC (Stewart et al., 2023), could benefit from BO’s sample efficiency, saving computational resources and time. Similarly, these benefits could transfer to existing frameworks for stellarator coil design (Jorge et al., 2024; Giuliani, 2024; Kaptanoglu et al., 2025).
1.3 Outline
@@
In Section 2, we describe the multi-objective BO problem, the Gaussian process surrogate model, the classifier scheme, and the acquisition function required in the BO loop. We follow this in Section 3 by defining the PF coil design problem in terms of the input space, the objectives we seek to optimise, and the constraints on the plasma and the machine. In addition, we present the simulator used to generate the plasma equilibria for each PF coil set and define cases in which the simulator may fail to produce a valid equilibrium (requiring the classifier). The numerical experiments are detailed and presented in Section 4. To highlight the data efficiency of the BO scheme, we carry out a number of experiments with a fixed computational budget and assess performance against alternative optimisation methods. In Section 5, we discuss our findings, highlight any major advantages and disadvantages of the BO framework applied to this problem, and propose avenues for future work.
2 Multi-objective Bayesian optimisation
@@
BO is a method for performing gradient-free global optimisation of black-box functions, typically utilised when the function is expensive-to-evaluate (Garnett, 2023). This is because practitioners will often want to identify (feasible) optimal points of the function’s input/output spaces with as few function evaluations as possible—especially if there is a limited computational budget.
@@
Here, we are interested in optimising the nonlinear function $\bm{f}\colon\mathcal{U}\subseteq\mathbb{R}^{d}\to\mathbb{R}^{l+m}$ that takes in a $d$-dimensional input and returns $l$ objectives and $m$ constraints. More formally, the aim of multi-objective BO ($l>1$), is to solve
@@
$\operatorname*{argmin}_{\begin{subarray}{c}\bm{x}\in\mathcal{U}\\ \bm{f}_{[l+1..l+m]}\preccurlyeq 0\end{subarray}}\bm{f}_{[1..l]}\left(\bm{x}\right),$ (2.1)
@@
Given we need to optimise over multiple competing objectives, problems such as (2.1) will often involve trade-offs where improving one objective may come at the expense of another. The aim is therefore to seek the set of Pareto optimal solutions $\mathcal{P}$ that are not dominated by any other solutions. A solution $\bm{x}$ dominates another $\bm{x}^{\prime}$, denoted $\bm{x}\prec\bm{x}^{\prime}$, if and only if $\bm{f}(\bm{x})\preccurlyeq\bm{f}(\bm{x}^{\prime})$ and $\exists j\in\{1,\ldots,l\}$ such that $\bm{f}_{j}(\bm{x})<\bm{f}_{j}(\bm{x}^{\prime})$. In short, a solution $\bm{x}$ dominates $\bm{x}^{\prime}$ if it is at least as good in all objectives and strictly better in at least one. Given a dataset
$\mathcal{D}=\left\{\left(\bm{x}_{i},\bm{f}(\bm{x}_{i})\right)\right\}_{i=1}^{N},$
consisting of $N$ evaluations of $\bm{f}$, the Pareto set for (2.1) is defined as
$\mathcal{P}(\mathcal{D})\coloneqq\{\bm{x}\in D\mid\nexists\ \bm{x^{\prime}}\in D\ \text{s.t.}\ \bm{x}^{\prime}\prec\bm{x}\}.$ (2.2)
@@
The Pareto front, denoted $\mathcal{P}_{\bm{f}}$, is defined as the image of the Pareto set, i.e. $\mathcal{P}_{\bm{f}}\coloneqq\{\bm{f}(\bm{x})\mid\bm{x}\in\mathcal{P}(\mathcal{D})\}$. See Garnett (2023, Chp. 11.7) for an illustration of the Pareto front.
2.1 The Bayesian optimisation loop
@@
The key component in BO for identifying feasible and optimal trade-offs between the objectives is a probabilistic surrogate model, capable of performing uncertainty based exploration. This model is typically trained on some initial dataset by maximising its marginal likelihood—more details on this surrogate model are given in Section 2.2.
@@
The first stage in BO (refer to Figure 2.1) is to construct this initial dataset (which we will call $\mathcal{D}$) by taking $N$ samples $\bm{x}\in\mathcal{U}$ and evaluating them all using $\bm{f}$. One popular method used is Sobol sampling (Sobol, 1967), whereby samples are chosen quasi-randomly with low discrepancy to achieve approximately uniform coverage of the input space. The number of samples $N$ chosen/required may depend on the size of $d$, the computational budget available, and if parallel processing is available (for the $\bm{f}$ evaluations). Note that at this point, while we could use $D$ to immediately generate a Pareto set $\mathcal{P}(\mathcal{D})$, this would almost certainly be a poor estimate given a lack of data points and that most would reside in non-optimal regions of the objective space.
@@
Figure 2.1: A flowchart illustrating the standard “BO loop” along with an additional failure region classifier step—see text for more details.
@@
It is worth noting that for many black-box functions such as $\bm{f}$, there will be failure regions of the input space that cannot be evaluated. The reasons for failure in our particular setting are discussed further in Section 3.2.5. During construction of the initial dataset, samples that lie in failure regions may be encountered and so we do not wish to include these in the dataset. We do, however, wish to learn from these samples so that we do not encounter similar samples again and therefore we train a classifier to predict when this may happen—a similar approach was taken by Hornsby et al. (2024) when generating gyrokinetic simulation datasets for spherical tokamaks. This classifier will be used within the BO loop, which can be seen in Figure 2.1 and is now outlined:
Stage 1:
@@
Generate the initial dataset $\mathcal{D}$ using the Sobol sampling scheme.
Stage 2:
Train the surrogate model using the dataset $\mathcal{D}$ (see Section 2.2) to obtain a probabilistic model: $\bm{f}(\bm{x})\mid\mathcal{D}$.
Stage 3:
@@
Train the failure region classifier using the data and the surrogate (see Section 2.3).
Stage 4:
@@
Maximise the acquisition function over the input space to identify the most “informative” point $\bm{x}^{*}$ to observe next, given the current surrogate model (and classifier) $\bm{f}(\bm{x}^{*})\mid\mathcal{D}$ (see Section 2.4).
Stage 5:
Evaluate $\bm{f}(\bm{x}^{*})$ and add it to the dataset: $\mathcal{D}\coloneqq\mathcal{D}\cup\{(\bm{x}^{*},\bm{f}(\bm{x}^{*}))\}$.
Stage 6:
@@
Check whether the terminating condition is met and if not, return to Stage 2. In our case, we check whether or not the maximum number of iterations $j$ has been exceeded (to remain within our computational budget) [^1] Other terminating conditions include stopping once improvements in the acquisition function are below some threshold or when the objectives are deemed to be sufficiently optimal (Ishibashi et al., 2023)..
2.2 The Gaussian process surrogate

### v2

@@ Abstract
The tokamak is a world-leading concept for producing sustainable energy via magnetically-confined nuclear fusion. Identifying where to position the magnets within a tokamak, specifically the poloidal field (PF) coils, is a design problem which requires balancing a number of competing economical, physical, and engineering objectives and constraints. In this paper, we show that multi-objective Bayesian optimisation (BO), an iterative optimisation technique utilising probabilistic machine learning models, can effectively explore this complex design space and return several optimal PF coilsets. These solutions span the Pareto front, a subset of the objective space that optimally satisfies the specified objective functions. We outline an easy-to-use BO framework and demonstrate that it outperforms alternative optimisation techniques while using significantly fewer computational resources. Our results show that BO is a promising technique for fusion design problems that rely on computationally demanding high-fidelity simulations.
I Introduction
I.1 Motivation and aims
A spherical tokamak is a torus-shaped device with a low aspect ratio that uses strong magnetic fields to confine and control a thermonuclear fusion plasma, with the goal of producing fusion energy (Freidberg, 2008). The Spherical Tokamak for Energy Production (STEP), currently in the design phase and targeting completion in 2040 (Chapman, Cowley, and Wilson, 2024), is one of a few ongoing fusion research and development projects based on the spherical tokamak concept. To deliver fusion power to the grid on such a short timescale, researchers are increasingly designing next-generation tokamaks in silico with multi-physics simulations, many of which require high-performance computing (HPC) resources.
@@ I.1 Motivation and aims
Our focus here will be on the design of the poloidal field (PF) coil system, which plays a critical role in controlling the position and shape of the plasma in both the core and divertor regions of the tokamak (Lim et al., 2010). In particular, some coils are crucial for managing the vertical stability of elongated plasmas, such as those in spherical tokamaks, where the higher elongation can lead to larger vertical instability, risking disruption without appropriate control (Anand et al., 2023). By generating poloidal magnetic fields, the PF coil system ensures the plasma remains in equilibrium, balancing the inward-facing magnetic forces produced by the coils against the outward-facing pressure-driven forces generated by the plasma (Wesson and Campbell, 2011). The design of the system—in terms of the coil positions, sizes, and shapes—will have a significant impact on plasma performance and stability and will therefore need to satisfy a number of competing (and often conflicting) constraints. In the plasma, for example, constraints are required to ensure X-points form in specific locations (for stability), strikepoints hit the correct divertor plates (for heat management), and total current density limits on the PF coils are not exceeded. In terms of the tokamak itself, the locations/sizes of the coils will inevitably be constrained by the vacuum vessel, diagnostic systems, and maintenance ports (to name but a few).
In addition to constraints, there will be a number of objectives related to the desired operational plasma conditions that we wish the chosen coilset to minimise or maximise (depending on the objective). This could include minimising the coil size to reduce fabrication, construction, and installation costs or could include minimising current flows to reduce power consumption and structural stresses from forces produced by the coils (Coleman and McIntosh, 2020). Moreover, we may wish to optimise certain properties of the plasma in the divertor chambers in order to minimise heat loads on plasma facing components and improve exhaust performance (Hudoba et al., 2023a, 2024). Simultaneously satisfying both the objectives and constraints will require the solution of a complex optimisation problem that needs to be tackled in a systematic, computationally efficient manner.
In this paper, we will perform multi-objective Bayesian optimisation (BO) on an earlier baseline design of the STEP PF coil system (Hudoba et al., 2023b). Our aims are to:
@@
    motivate more widespread adoption of BO for the in silico design of interlinked components on future tokamak devices to save time, minimise financial costs, and improve plasma performance.
@@ I.2 Related work
Despite these advantages, the adoption and application of BO in fusion engineering and design has, so far, remained relatively limited. Brown et al. (2024)(Brown, Cioba, and Bogunovic, 2024) aimed to improve six key properties of the safety factor profile by using BO on the current profiles in STEP. They also demonstrate that BO performs better than a genetic algorithm with the same number of black-box function evaluations (as we will do later on). Mehta et al. (2024) use BO to find the parameters such as neutral beam injection power, plasma current, and plasma elongation in the DIII-D tokamak that safeguard against disruption during the ramp-down phase. Similarly, Pusztai et al. (2023) use BO to mitigate the impact of disruptions in ITER by exploring how injected deuterium and neon can minimise runaway electron currents, transported heat, and quench time post-disruption. Järvinen et al. (2022) also investigate runaway electron currents using BO as an advanced sampling method to help calibrate uncertainty and minimise the discrepancy between simulations and experimental data. For fusion component design, Humphrey et al. (2023) demonstrate the use of BO to minimise stresses in parametrised divertor monoblocks under fusion conditions. The most relevant work to ours is that of Nunn, Gopakumar, and Kahn (2023), who use multi-objective BO to optimise TF coil shapes to reduce both financial costs and magnetic ripples (which affect plasma stability and performance). In contrast, our approach deals with more computationally expensive, failure-prone plasma equilibrium simulations without analytic objective/constraint functions, necessitating the use of a classifier alongside the surrogate model.
@@
The work here is inspired by that of Hudoba et al. (2023b), in which the authors seek to optimise the STEP PF coil system by minimising deviations of key plasma parameters from a baseline scenario (which we adopt) and coil currents, while maximising divertor performance metrics. Using a free-boundary equilibrium solver, thousands of potential PF coilsets are sampled and evaluated (in a Monte Carlo-type approach) before optimal solution sets are identified heuristically. We aim to provide and fully outline an alternative, much more data efficient, framework for carrying out similar multi-objective optimisation that can return a STEP equilibrium similar to the baseline.
@@
There are also a number of areas in fusion design where BO has yet to be applied but could potentially offer significant benefits. For example, parameter scans for optimal magnetic sensor placement, as explored for TCV (Romero and Svensson, 2013) and SPARC (Stewart et al., 2023), could benefit from BO’s sample efficiency, saving computational resources and time. Similarly, these benefits could transfer to existing frameworks for stellarator coil design (Jorge, Giuliani, and Loizu, 2024; Giuliani, 2024; Kaptanoglu, Landreman, and Zarnstorff, 2025).
I.3 Outline
@@
In Section II, we describe the multi-objective BO problem, the Gaussian process surrogate model, the classifier scheme, and the acquisition function required in the BO loop. We follow this in Section III by defining the PF coil design problem in terms of the input space, the objectives we seek to optimise, and the constraints on the plasma and the machine. In addition, we describe the simulator used to generate the plasma equilibria for each PF coilset and define cases in which the simulator may fail to produce a valid equilibrium (requiring the classifier). The numerical experiments are detailed and presented in Section IV. To highlight the data efficiency of the BO scheme, we carry out a number of experiments with a fixed computational budget and assess performance against alternative optimisation methods. In Section V, we discuss our findings, highlight any major advantages and disadvantages of the BO framework applied to this problem, and propose avenues for future work.
II Multi-objective Bayesian optimisation
@@
BO is a method for performing gradient-free global optimisation of black-box functions, typically utilised when the function is expensive-to-evaluate (Garnett, 2023). Practitioners will often want to identify (feasible) optimal points of the function’s input/output spaces with as few function evaluations as possible—especially if there is a limited computational budget.
@@
Here, we are interested in optimising the nonlinear function $\bm{f}\colon\mathscr{U}\subseteq\mathbb{R}^{d}\to\mathbb{R}^{l+m}$ that takes in a $d$-dimensional input and returns $l$ objectives and $m$ constraints. More formally, the aim of multi-objective BO ($l>1$) is to solve
@@
$\operatorname*{argmin}_{\begin{subarray}{c}\bm{x}\in\mathscr{U}\\ \bm{f}_{[l+1..l+m]}\preccurlyeq 0\end{subarray}}\bm{f}_{[1..l]}\left(\bm{x}\right),$ (1)
@@
Given we need to optimise over multiple competing objectives, problems such as (1) will often involve trade-offs where improving one objective may come at the expense of another. The aim is therefore to seek the set of Pareto optimal solutions $\mathscr{P}$ that are not dominated by any other solutions. A solution $\bm{x}$ dominates another $\bm{x}^{\prime}$, denoted $\bm{x}\prec\bm{x}^{\prime}$, if and only if $\bm{f}(\bm{x})\preccurlyeq\bm{f}(\bm{x}^{\prime})$ and $\exists j\in\{1,\ldots,l\}$ such that $\bm{f}_{j}(\bm{x})<\bm{f}_{j}(\bm{x}^{\prime})$. In short, a solution $\bm{x}$ dominates $\bm{x}^{\prime}$ if it is at least as good in all objectives and strictly better in at least one. Given a dataset
$\mathscr{D}=\left\{\left(\bm{x}_{i},\bm{f}(\bm{x}_{i})\right)\right\}_{i=1}^{N},$
consisting of $N$ evaluations of $\bm{f}$, the Pareto set for (1) is defined as
$\mathscr{P}(\mathscr{D})\coloneqq\{\bm{x}\in\mathscr{D}\mid\nexists\ \bm{x^{\prime}}\in\mathscr{D}\ \text{s.t.}\ \bm{x}^{\prime}\prec\bm{x}\}.$ (2)
@@
The Pareto front, denoted $\mathscr{P}_{\bm{f}}$, is defined as the image of the Pareto set, i.e. $\mathscr{P}_{\bm{f}}\coloneqq\{\bm{f}(\bm{x})\mid\bm{x}\in\mathscr{P}(\mathscr{D})\}$. See Garnett (2023, Chp. 11.7) for an illustration of the Pareto front.
II.1 The Bayesian optimisation loop
@@
The key component in BO for identifying feasible and optimal trade-offs between the objectives is a probabilistic surrogate model, capable of performing uncertainty based exploration. This model is typically trained on some initial dataset by maximising its marginal likelihood—more details on this surrogate model are given in Section II.2.
@@
The first stage in BO (refer to Figure 1) is to construct this initial dataset (which we will call $\mathscr{D}$) by taking $N$ samples $\bm{x}\in\mathscr{U}$ and evaluating them all using $\bm{f}$. One popular method used is Sobol sampling (Sobol, 1967), whereby samples are chosen quasi-randomly with low discrepancy to achieve approximately uniform coverage of the input space. The number of samples $N$ chosen/required may depend on the size of $d$, the computational budget available, and if parallel processing is available (for the $\bm{f}$ evaluations). Note that at this point, while we could use $\mathscr{D}$ to immediately generate a Pareto set $\mathscr{P}(\mathscr{D})$, this would almost certainly be a poor estimate given a lack of data points and that most would reside in non-optimal regions of the objective space.
@@
Figure 1: A flowchart illustrating the standard “BO loop” along with an additional failure region classifier step—see text for more details.
@@
It is worth noting that for many black-box functions such as $\bm{f}$, there will be failure regions of the input space that cannot be evaluated. The reasons for failure in our particular setting are discussed further in Section III.2.5. During construction of the initial dataset, samples that lie in failure regions may be encountered and so we do not wish to include these in the dataset. We do, however, wish to learn from these samples so that we do not encounter similar samples again and therefore we train a classifier to predict when this may happen—a similar approach was taken by Hornsby et al. (2024) when generating gyrokinetic simulation datasets for spherical tokamaks. This classifier will be used within the BO loop, which can be seen in Figure 1 and is now outlined:
Stage 1
@@
Generate the initial dataset $\mathscr{D}$ using the Sobol sampling scheme.
Stage 2
Train the surrogate model using the dataset $\mathscr{D}$ (see Section II.2) to obtain a probabilistic model: $\bm{f}(\bm{x})\mid\mathscr{D}$. Note that we exclude failure samples during training as they do not provide any valid information about the objective or constraint values.
Stage 3
@@
Train the failure region classifier using the generated data (see Section II.3).
Stage 4
@@
Maximise the acquisition function over the input space to identify the most “informative” point $\bm{x}^{*}$ to observe next, given the current surrogate model (and classifier) $\bm{f}(\bm{x}^{*})\mid\mathscr{D}$ (see Section II.4).
Stage 5
Evaluate $\bm{f}(\bm{x}^{*})$ and add it to the dataset: $\mathscr{D}\coloneqq\mathscr{D}\cup\{(\bm{x}^{*},\bm{f}(\bm{x}^{*}))\}$.
Stage 6
@@
Check whether the terminating condition is met and if not, return to Stage 2. In our case, we check whether or not the maximum number of iterations has been exceeded (to remain within our computational budget). Other terminating conditions include stopping once improvements in the acquisition function are below some threshold or when the objectives are deemed to be sufficiently optimal (Ishibashi et al., 2023).
II.2 The Gaussian process surrogate